An airplane accelerates down a runway at 3.20 m/s2 for 32.8 s until is finally lifts off the ground. Determine the distance traveled before takeoff.
1
Expert's answer
2020-12-23T10:25:13-0500

The distance traveled under the constant acceleration "a = 3.2m\/s^2" in time "t = 32.8s" is (assume starting from rest):


"d = \dfrac{at^2}{2}\\
d = \dfrac{3.2\cdot 32.8^2}{2} \approx 1721m"

Answer. 1721 m.